# Enhancing Low-light Light Field Images with A Deep Compensation Unfolding Network

### v2

@@ V-C Ablation Study
TABLE IV: Quantitative results of ablation studies assessing the impact of various network architecture components within the proposed method. These components include network modules such as Illumination Map Estimation (IE), Content-Associated Deep Compensation (CDC), and Deep Optimization (DO), as well as DPER structures comprising the Simplified Module (SIM), EPI Boost Structure ($f_{boost}$), and Ray Mixing Structure ($f_{mix}$). $\checkmark$” (resp. “$\times$”) represents the corresponding structure is used (resp. unused).
@@
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.56 0.878 0.108
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.21 0.871 0.116
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 33.50 0.879 0.110
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 33.18 0.872 0.114
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 33.62 0.884 0.104
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 33.67 0.885 0.104
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.96 0.886 0.103
@@
The pseudo-explicit feature interaction is designed for effective feature extraction of low-light LFs without depth information by converging rays and boosting EPI convolution. To validate the efficacy of the proposed structures $f_{boost}$ and $f_{mix}$, we incrementally introduced them into the simplified module (SIM) base configuration. The ablation study results, as delineated in Table IV, reveal that both $f_{boost}$ and $f_{mix}$ contribute uniquely to the feature extraction process. It is evident that the full deployment of the proposed DPEF achieves the best performance.
To demonstrate the superiority of our proposed pseudo-explicit feature interaction, we performed comparative experiments by replacing our module with Spatial-Angular Separable convolution [42] and Disentangling convolution [51]. For clarity, we refer to these variants as Ours-SAS and Ours-Distg, respectively. We adjusted the number of feature channels to align the overall parameter count across models for a fair comparison. The results presented in Table V indicate that our DCUNet, incorporating the proposed pseudo-explicit feature interaction with specialized techniques such as EPI boost convolution and ray convergence, tailored for low-light LF enhancement, achieves superior performance over the Ours-SAS and Ours-Distg models. This clearly validates the benefits of our pseudo-explicit feature interaction approach.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours-SAS 32.38 0.853 0.121
Ours-Distg 32.74 0.860 0.129
Ours 33.96 0.886 0.103
@@
Fixed Dynamic PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
$S=3$ $L=2$ 33.02 0.866 0.117
$L=4$ 33.59 0.878 0.105
$L=6$ 33.96 0.886 0.103
$L=8$ 34.02 0.890 0.102
$L=6$ $S=1$ 33.34 0.873 0.111
$S=2$ 33.67 0.881 0.105
$S=3$ 33.96 0.886 0.103
$S=4$ 34.15 0.890 0.101
@@
Loss configurations. We investigated the effectiveness of different loss components, namely L1 loss $L_{1}$, SSIM loss $\texttt{SSIM}(\cdot)$, and perceptual loss $\texttt{Per}(\cdot)$. The results in Table VIII demonstrate a gradual improvement in performance, indicating that incorporating multiple loss components could enhance our DCUNet’s performance.

### v3

@@ V-C Ablation Study
TABLE IV: Quantitative results of ablation studies assessing the impact of various network architecture components within DCUNet. These components include network modules such as Illumination Map Estimation (IE), Content-Associated Deep Compensation (CDC), and Deep Optimization (DO), as well as DPER structures comprising the Simplified Module (SIM), EPI Boost Structure ($f_{boost}$), and Ray Mixing Structure ($f_{mix}$). $\checkmark$” (resp. “$\times$”) represents the corresponding structure is used (resp. unused).
@@
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.56 0.878 0.108 3.84
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.21 0.871 0.116 3.83
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 33.50 0.879 0.110 3.85
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 33.37 0.875 0.112 3.95
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 33.64 0.884 0.104 3.86
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 33.72 0.884 0.104 3.90
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.96 0.886 0.103 3.84
@@
The pseudo-explicit feature interaction is designed for effective feature extraction of low-light LFs without depth information by converging rays and boosting EPI convolution. To validate the efficacy of the proposed structures $f_{boost}$ and $f_{mix}$, we incrementally introduced them into the simplified module (SIM) base configuration and adjusted feature channel numbers to align the overall parameter. The ablation study results, as delineated in Table IV, reveal that both $f_{boost}$ and $f_{mix}$ contribute uniquely to the feature extraction process. It is evident that the full deployment of the proposed DPEF achieves the best performance.
To demonstrate the superiority of our proposed pseudo-explicit feature interaction, we performed comparative experiments by replacing our module with Spatial-Angular Separable convolution [44], Spatial-Angular Versatile (SAV) convolution [47] and Disentangling convolution [46]. For clarity, we refer to these variants as Ours-SAS, Ours-SAV and Ours-Distg, respectively. We adjusted the number of feature channels to align the overall parameter count across models for a fair comparison. The strategy of full-angular convolution combined with $1\times 1$ convolution to generate all angular information in Disentangling convolution may not be suitable for processing low-light light field images with large baseline and severe noise. The results presented in Table V indicate that our DCUNet, incorporating the proposed pseudo-explicit feature interaction with specialized techniques such as EPI boost convolution and ray convergence, tailored for low-light LF enhancement, achieves superior performance over the Ours-SAS, Ours-SAV, and Ours-Distg models. This clearly validates the benefits of our pseudo-explicit feature interaction approach.
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ #Params(M)
Ours-SAS 32.38 0.853 0.121 4.02
Ours-SAV 32.09 0.852 0.135 4.05
Ours-Distg 32.74 0.860 0.129 3.62
Ours 33.96 0.886 0.103 3.84
@@
Fixed Dynamic PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ #Params(M)
$S=3$ $L=2$ 33.02 0.866 0.117 1.26
$L=4$ 33.59 0.878 0.105 2.54
$L=6$ 33.96 0.886 0.103 3.84
$L=8$ 34.02 0.890 0.102 5.20
$L=6$ $S=1$ 33.34 0.873 0.111 1.28
$S=2$ 33.67 0.881 0.105 2.57
$S=3$ 33.96 0.886 0.103 3.84
$S=4$ 34.15 0.890 0.101 5.14
@@
Loss configurations. We investigated the effectiveness of different loss components, namely L1 loss $L_{1}$, SSIM loss $\texttt{SSIM}(\cdot)$, and perceptual loss $\texttt{Per}(\cdot)$. The results in Table VIII demonstrate a gradual performance improvement, indicating that incorporating multiple loss components could enhance our DCUNet’s performance.